# Cross-View Geo-Localization with Street-View and VHR Satellite Imagery in Decentrality Settings

### v1

@@ 6.3 Visualization Analysis
To qualitatively evaluate AuxGeo’s effectiveness, we visualize heatmaps of features from the baseline and AuxGeo models using images from the test set of CVSat. As shown in Figure 8, the upper two rows depict a narrow neighborhood scene, while the lower two rows depict an open crossroads scene. The heatmaps reveal that, compared to the baseline (Sample4Geo), AuxGeo more effectively focuses on salient features like roads, trees, and corners, which better emphasizes co-visible regions of cross-view images.
@@ 7 Conclusion
In this work, we introduce the concept of decentrality and propose CVSat, a novel dataset for cross-view geo-localization that allows for a thorough investigation of decentrality issues. This dataset comprises cross-view images from eight global cities, representing diverse styles and encompassing extensive regions, including urban, suburban, and rural areas. To tackle both the cross-view problem and the issue of decentrality, we propose a novel method called AuxGeo. AuxGeo employs a multi-metric optimization strategy incorporating two novel auxiliary modules designed to enhance the backbone’s representational ability. Extensive experiments demonstrate that AuxGeo achieves state-of-the-art performance on our proposed CVSat as well as on public datasets such as CVUSA, CVACT, and VIGOR. Importantly, AuxGeo incurs no additional pre-processes or computational costs during inference. While AuxGeo achieves a favorable balance between localization accuracy and efficiency, it falls short in addressing the low accuracy of CVGL under conditions of extremely high decentrality. This limitation highlights an important avenue for future research.
@@ Acknowledgment
This work was supported by the National Natural Science Foundation of China (Grants 42471470, 42030102, and 42192583) and Ant Group.

### v2

@@ 6.3 Visualization Analysis
To qualitatively evaluate AuxGeo’s effectiveness, we visualize heatmaps of features from the baseline and AuxGeo models using images from the test set of DReSS. As shown in Figure 8, the upper two rows depict a narrow neighborhood scene, while the lower two rows depict an open crossroads scene. The heatmaps reveal that, compared to the baseline (Sample4Geo), AuxGeo more effectively focuses on salient features like roads, trees, and corners, which better emphasizes co-visible regions of cross-view images.
@@ 7 Conclusion
In this work, we introduce the concept of decentrality and propose DReSS, a novel dataset for cross-view geo-localization that allows for a thorough investigation of decentrality issues. This dataset comprises cross-view images from eight global cities, representing diverse styles and encompassing extensive regions, including urban, suburban, and rural areas. To tackle both the cross-view problem and the issue of decentrality, we propose a novel method called AuxGeo. AuxGeo employs a multi-metric optimization strategy incorporating two novel auxiliary modules designed to enhance the backbone’s representational ability. Extensive experiments demonstrate that AuxGeo achieves state-of-the-art performance on our proposed DReSS as well as on public datasets such as CVUSA, CVACT, and VIGOR. Importantly, AuxGeo incurs no additional pre-processes or computational costs during inference. While AuxGeo achieves a favorable balance between localization accuracy and efficiency, it falls short in addressing the low accuracy of CVGL under conditions of extremely high decentrality. This limitation highlights an important avenue for future research.
@@ Acknowledgment
This work was supported by the National Natural Science Foundation of China (Grants 42471470, 42030102, and 42192583), Ant Group and the Supercomputing Center of Wuhan University.
[…]